If you add​ Natalie's age and​ Fred's age, the result is 38. If you add​ Fred's age to 4 times​ Natalie's age, the result is 74. Write and solve a system of equations to find how old Fred and Natalie are.
Found 2 solutions by jim_thompson5910, greenestamps:
Answer by jim_thompson5910(35256)


N = Natalie's age
F = Fred's age
both ages are in years, and they are current present day ages

"If you add Natalie's age and Fred's age, the result is 38." means the first equation is F+N = 38.
Solve for F to get F = 38 - N. You subtract N from both sides. We'll use this equation later.

"If you add Fred's age to 4 times Natalie's age, the result is 74." so the second equation is
F + 4*N = 74

Now plug in F = 38 - N. Solve for N.
F + 4*N = 74
38 - N + 4*N = 74 ... replaced F with 38-N
38 + 3N = 74
38 + 3N-38 = 74-38 .... subtracting 38 from both sides
3N = 36
3N/3 = 36/3 .... divide both sides by 3
N = 12

Use this to find Fred's age
F = 38 - N
F = 38 - 12 ... plug in N = 12
F = 26

-------------------------------------------------
To summarize, the system of equations is
F + N = 38
F + 4N = 74

Solving said system leads to N = 12 and F = 26. This means Natalie is 12 years old and Fred is 26 years old.

Answer by greenestamps(13200)



The two equations, as shown by the other tutor, are indeed

F+N = 38
F+4N = 74

But with the two equations in that form, using substitution is the long way around; the two equations are perfectly suited for solving by elimination.

Subtracting the first equation from the second yields

3N = 36

which immediately gives us

N = 12

and then the first equation gives us

F = 26